# Active Sensing for Reciprocal MIMO Channels

### v1

@@ Abstract
This paper addresses the design of transmit precoder and receive combiner matrices to support $N_{\rm s}$ independent data streams over a time-division duplex (TDD) point-to-point massive multiple-input multiple-output (MIMO) channel with either a fully digital or a hybrid structure. The optimal precoder and combiner design amounts to finding the top-$N_{\rm s}$ singular vectors of the channel matrix, but the explicit estimation of the entire high-dimensional channel would require significant pilot overhead. Alternatively, prior works seek to find the precoding and combining matrices directly by exploiting channel reciprocity and by using the power iteration method, but its performance degrades in the low SNR regime. To tackle this challenging problem, this paper proposes a learning-based active sensing framework, where the transmitter and the receiver send pilots alternately using sensing beamformers that are actively designed as functions of previously received pilots. This is accomplished by using recurrent neural networks to summarize information from the historical observations into hidden state vectors, then using fully connected neural networks to learn the appropriate sensing beamformers in the next pilot stage and finally the transmit precoding and receive combiner matrices for data communications. Simulations demonstrate that the learning-based method outperforms existing approaches significantly and maintains superior performance even in low SNR regimes both in fully digital and hybrid MIMO scenarios.
@@ I Introduction
The conventional approach to massive MIMO system design involves first estimating the channel and then optimizing the beamformers based on the estimated channel [8, 9]. To reduce the pilot training overhead, the knowledge of the underlying structure of the channel, e.g., sparsity in mmWave channels, can be utilized in the channel estimation step [10, 11, 12]. However, such an approach relies heavily on the geometry of the antenna arrays and the sparsity assumption on the channel model, which is not necessarily accurate. Moreover, in conventional approaches, the sensing beamformers for pilot transmission are often generated randomly, but sensing in random directions is not necessarily optimal, because typically only a low-dimensional subspace of the channel is of interest for designing the precoder and the beamformers.
Inspired by recent developments in deep learning for reducing pilot training overhead [13], this paper proposes a learning-based sensing framework that directly learns the beamformers from received pilots without estimating the entire channel and in an active fashion over multiple rounds. The main idea is to utilize a set of carefully designed recurrent neural networks (RNNs) and fully connected deep neural networks (DNNs) to automatically summarize information from the historical observations into state vectors, which are used to design the sensing beamformers in the next stage, so that the sensing beamformers gradually focus the energy towards to directions of interest, e.g., the directions of the top-$N_{\rm s}$ singular vectors. We show that the proposed framework can work well in both fully digital and hybrid beamforming scenarios and is robust to the noise.
@@ II-A System Model
This paper addresses the case where the channel matrix $\bm{G}$ is unknown. A conventional technique for designing $\bm{W}_{\rm t}$ and $\bm{W}_{\rm r}$ would have required to first estimate the matrix channel in a pilot stage, followed by the SVD of the estimated channel. However, the estimating the entire channel matrix would require significant pilot overhead, especially for massive MIMO systems.
To reduce pilot training overhead, we make a key observation that the optimal transmit precoding and receive combiner matrices, i.e., (II-A), are only functions of the top-$N_{\rm s}$ singular vectors of $\bm{G}$, rather than the entire channel matrix. Thus, it should be possible to directly learn the transmit precoding and receive combiner matrices by probing only low-dimensional subspaces of the channel matrix.
@@ II-B Ping-Pong Pilots Transmission
$\bm{W}_{\ell}^{\rm A}=f_{\ell}^{\rm A}([\bm{Y}_{0}^{\rm A},\ldots,\bm{Y}_{\ell-1}^{\rm A}]),\quad\ell=1,\ldots,L-1,$ (12a)
@@
$\underset{\mathcal{F},\bm{Q}_{\rm v},\bm{Q}_{\rm u}}{\operatorname{minimize}}~{}$ $\mathbb{E}\left[\|{\bm{W}}_{{\rm t}}-\bm{V}_{1}\bm{Q}_{\rm v}\|_{2}^{2}+\|{\bm{W}}_{{\rm r}}-\bm{U}_{1}\bm{Q}_{\rm u}\|_{2}^{2}\right]$ (14)
@@
Preparing a sufficient amount of training data is crucial in the data-driven approach. However, the objective function defined in (14) is not easy to deal with, because optimization over $\bm{Q}_{\rm v}$ and $\bm{Q}_{\rm u}$ is involved, and further it requires performing SVD for every channel matrix in the training data set in order to obtain $\bm{V}_{1}$ and $\bm{U}_{1}$. To avoid this complexity, we propose to replace the objective function in (14) by $-\mathbb{E}\left[\log\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})\right]$, because the function $\log\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})$ is maximized when ${\bm{W}}_{{\rm t}}=\bm{V}_{1}$ and ${\bm{W}}_{{\rm r}}=\bm{U}_{1}$. In this way, we reformulate the active sensing problem as:
$\underset{\mathcal{F}}{\operatorname{maximize}}~{}$ $\mathbb{E}\left[\log\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})\right]$ (15)
@@ III-A Motivations From Power Iteration Method
where we use the fact that $\bm{w}_{0}^{\rm A}$ can be expressed as $\bm{w}_{0}^{\rm A}=\sum_{i=1}^{\operatorname{rank}(\bm{G})}\beta_{i}\bm{v}_{i}+\bar{\bm{v}}_{{\rm t},0}$, with $\bar{\bm{v}}_{{\rm t},0}$ belonging to the null space of $\bm{G}$. As the number of transmission rounds $\ell$ increase, the vectors $\bm{y}_{\ell}^{\rm A}$ and $\bm{y}_{\ell}^{\rm B}$ in (III-A) would be dominated by the top singular vectors $\bm{v}_{1}$ and $\bm{u}_{1}$, respectively, with a linear convergence rate [7]. Furthermore, to meet the transmit power constraint, the unit 2-norm normalization is performed in each iteration, i.e., $\bm{w}_{\ell}^{\rm B}={\bm{y}_{\ell}^{\rm B}}/{\|\bm{y}_{\ell}^{\rm B}\|_{2}}$ and $\bm{w}_{\ell+1}^{\rm A}={\bm{y}_{\ell}^{\rm A}}/{\|\bm{y}_{\ell}^{\rm A}\|_{2}}$, which does not alter the directions of $\bm{y}_{\ell}^{\rm A}$ and $\bm{y}_{\ell}^{\rm B}$. The rapid convergence rate of the power iteration scheme implies a significant pilot overhead reduction as compared to estimating the entire channel matrix.
@@ III-B Proposed Active Sensing Approach
Similarly, agent A uses its hidden state vector $\bm{h}_{L,i}^{\rm A}$ and received pilots $\bm{Y}_{L-1}^{\rm A}$ to design its precoding matrix $\bm{W}_{\rm t}$ following the same procedure as in (III-B), but replaces the $\operatorname{DNN}^{\rm r}$ and $\bm{Y}_{L-1}^{\rm B}$ with $\operatorname{DNN}^{\rm t}$ and $\bm{Y}_{L-1}^{\rm A}$, respectively.
The proposed active sensing framework is trained by sequentially concatenating $L$ active sensing units to learn the global policy for the entire sensing trajectory during the training stage. The loss function is set as $-\mathbb{E}\left[\log\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})\right]$. Furthermore, to train a neural network to be generalizable to different number of ping-pong rounds $L$, we can set the loss function as $-\mathbb{E}\left[\sum_{\ell=0}^{L-1}\log\det(\bm{W}_{{\rm r},\ell}^{\sf H}\bm{G}\bm{W}_{{\rm t},\ell})\right]$, where $\bm{W}_{{\rm r},\ell}$ and $\bm{W}_{{\rm t},\ell}$ are the data transmission precoding and combining matrices generated by the neural network in round $\ell$.
@@ IV Performance Evaluation for Fully Digital MIMO Systems
In this section, we evaluate the performance of the proposed active sensing framework for fully digital MIMO systems. The dimensions of the hidden states of the GRUs are set to be $N_{\rm h}^{\rm A}=N_{\rm h}^{\rm B}=512$. All the DNNs are two-layer fully connected neural networks with the dimension $[512,1024,2M_{\rm t}]$ and $\operatorname{relu}(\cdot)$ as the activation function in the hidden layers. The entire neural network involving $L$ concatenated active sensing units (as shown in Fig. 2) is implemented on PyTorch [30], which can automatically compute the gradient of the loss function with respect to the trainable parameters. We train the neural network offline using the Adam optimizer [31], with a default configuration in PyTorch and with an initial learning rate progressively decreasing from $10^{-3}$ to $10^{-5}$. In each training iteration, the batch size of training data is $1024$. We generate as much training data as needed and stop training if the performance on the validation dataset does not improve after several epochs. The testing dataset consists of $1000$ randomly generated channel matrices. In the training stage, we use the loss function $-\mathbb{E}\left[\sum_{\ell=0}^{L-1}\log\det(\bm{W}_{{\rm r},\ell}^{\sf H}\bm{G}\bm{W}_{{\rm t},\ell})\right]$ to train a single model that is generalizable to different number of ping-pong rounds.
@@
Figure 3: Performance comparison under Rayleigh fading channel model for a fully digital MIMO system with $M_{\rm t}=M_{\rm r}=64$.
@@
Figure 5: Performance comparison under ray-tracing channel model for a fully digital MIMO system with $M_{\rm t}=64,M_{\rm r}=16$ and $N_{\rm s}=2$.
@@ IV-A Rayleigh Fading Channel
In Fig. 3, we plot the performance comparison of different methods regarding the average objective value for varying SNRs and pilot lengths. The results of the scenario with $N_{\rm s}=2$ and $N_{\rm s}=4$ data streams are shown in Fig. 3a3d and Fig. 3e3h, respectively. In both scenarios, the neural network is trained for the scenario where the number of ping-pong rounds is $L=8$ (corresponding to pilot length $322LN_{\rm s}=$ 32 for $N_{\rm s}=2$ and $642LN_{\rm s}=$ 64 for $N_{\rm s}=4$).
@@
In Fig. 3e3h, we further plot the performance comparison of different methods for the scenario $N_{\rm s}=4$. It is observed that the proposed active sensing method still achieves the best performance. However, by comparing Fig. 3c and Fig. 3g, we can see that the conventional power iteration method and summed power method is not as competitive as the channel estimation based approach when the number of data streams increases. We should also note that the conventional channel estimation based approach requires an additional error-free feedback channel to transmit the designed precoding matrix from agent B to A. The feedback channel may not be easy to realize when the SNR is low. The other approaches do not need the feedback channel since the two agents can implicitly coordinate at the ping-pong pilot transmission stage.
In Fig. 4, we further present the theoretical achievable rate (without employing water-filling) with $N_{\rm s}=2$, $M_{\rm r}=64$ and $M_{\rm t}=64$, where the data transmission SNR is set at $10$dB. From Fig. 4, it is evident that the proposed active sensing method outperforms other approaches across various pilot SNRs. This figure shows that optimizing the proposed objective function $\log\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})$ also approximately maximizes the achievable rate.
@@ V-A System Model
Unlike the fully digital scenario, the problem of finding the optimal hybrid beamforming matrices does not have an analytical solution and is more challenging even when the channel matrix $\bm{G}$ is perfectly known. For the fully digital MIMO system, the optimal overall precoding matrix $\bm{W}_{\rm t}^{\ast}$ and overall combining matrix $\bm{W}_{\rm r}^{\ast}$ can be found as in (II-A), but it is not clear whether a further matrix decomposition step on $\bm{W}_{\rm t}^{\ast}$ and $\bm{W}_{\rm r}^{\ast}$ can be performed to obtain the optimal analog and digital beamformers. One approach for finding the hybrid precoding beamformers is to formulate the following matrix decomposition problem [33]:

### v2

@@ Abstract
This paper addresses the design of transmit precoder and receive combiner matrices to support $N_{\rm s}$ independent data streams over a time-division duplex (TDD) point-to-point massive multiple-input multiple-output (MIMO) channel with either a fully digital or a hybrid structure. The optimal precoder and combiner design amounts to finding the top-$N_{\rm s}$ singular vectors of the channel matrix, but the explicit estimation of the entire high-dimensional channel would require significant pilot overhead. Alternatively, prior works suggest to find the precoding and combining matrices directly by exploiting channel reciprocity and by using the power iteration method, but its performance degrades in the low SNR regime. To tackle this challenging problem, this paper proposes a learning-based active sensing framework, where the transmitter and the receiver send pilots alternately using sensing beamformers that are actively designed as functions of previously received pilots. This is accomplished by using recurrent neural networks to summarize information from the historical observations into hidden state vectors, then using fully connected neural networks to learn the appropriate sensing beamformers in the next pilot stage and finally the transmit precoding and receive combiner matrices for data communications. Simulations demonstrate that the learning-based method outperforms existing approaches significantly and maintains superior performance even in the low SNR regime for both the fully digital and hybrid MIMO scenarios.
@@ I Introduction
The conventional approach to massive MIMO system design involves first estimating the channel and then optimizing the beamformers based on the estimated channel [8, 9]. To reduce the pilot training overhead, the knowledge of the underlying structure of the channel, e.g., sparsity in mmWave channels, can be utilized in the channel estimation step [10, 11, 12]. However, such an approach relies heavily on the geometry of the antenna arrays and the sparsity assumption on the channel model, which is not necessarily accurate. Moreover, in conventional approaches, the sensing beamformers for pilot transmission are often generated randomly, but sensing in random directions is not necessarily optimal, because typically only a low-dimensional subspace of the channel is of interest for designing the precoder and the combiner.
Inspired by recent developments in deep learning for reducing pilot training overhead [13], this paper proposes a learning-based sensing framework that directly learns the beamformers from received pilots without estimating the entire channel and in an active fashion over multiple rounds. The main idea is to utilize a set of carefully designed recurrent neural networks (RNNs) and fully connected deep neural networks (DNNs) to automatically summarize information from the historical observations into state vectors, which are used to design the sensing beamformers in the next stage, so that the sensing beamformers gradually focus the energy towards the directions of interest, e.g., the directions of the top-$N_{\rm s}$ singular vectors. We show that the proposed framework can work well in both fully digital and hybrid beamforming scenarios and is robust to the noise.
@@ II-A System Model
This paper addresses the case where the channel matrix $\bm{G}$ is unknown. A conventional technique for designing $\bm{W}_{\rm t}$ and $\bm{W}_{\rm r}$ would have required to first estimate the matrix channel in a pilot stage, followed by the SVD of the estimated channel. However, estimating the entire channel matrix would require significant pilot overhead, especially in massive MIMO systems.
To reduce pilot training overhead, we make a key observation that the optimal transmit precoding and receive combiner matrices, i.e., (5), are only functions of the top-$N_{\rm s}$ singular vectors of $\bm{G}$, rather than the entire channel matrix. Thus, it should be possible to directly learn the transmit precoding and receive combiner matrices by probing only low-dimensional subspaces of the channel matrix.
@@ II-B Ping-Pong Pilots Transmission
$\bm{W}_{\ell+1}^{\rm A}=f_{\ell}^{\rm A}([\bm{Y}_{0}^{\rm A},\ldots,\bm{Y}_{\ell}^{\rm A}]),\quad\ell=0,\ldots,L-2,$ (8a)
@@
$\underset{\mathcal{F},\bm{Q}_{\rm v},\bm{Q}_{\rm u}}{\operatorname{minimize}}~{}$ $\mathbb{E}\left[\|{\bm{W}}_{{\rm t}}-\bm{V}_{1}\bm{Q}_{\rm v}\|_{F}^{2}+\|{\bm{W}}_{{\rm r}}-\bm{U}_{1}\bm{Q}_{\rm u}\|_{F}^{2}\right]$ (10)
@@
Preparing a sufficient amount of training data is crucial in the data-driven approach. However, the objective function defined in (10) is not easy to deal with, because optimization over $\bm{Q}_{\rm v}$ and $\bm{Q}_{\rm u}$ is involved, and further it requires performing SVD for every channel matrix in the training data set in order to obtain $\bm{V}_{1}$ and $\bm{U}_{1}$. To avoid this complexity, we propose to replace the objective function in (10) by $-\mathbb{E}\left[\log|\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})|^{2}\right]$, because the function $\log|\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})|^{2}$ is maximized when ${\bm{W}}_{{\rm t}}=\bm{V}_{1}$ and ${\bm{W}}_{{\rm r}}=\bm{U}_{1}$. In this way, we reformulate the active sensing problem as:
$\underset{\mathcal{F}}{\operatorname{maximize}}~{}$ $\mathbb{E}\left[\log|\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})|^{2}\right]$ (11)
@@ III-A Motivations From Power Iteration Method
where we use the fact that $\bm{w}_{0}^{\rm A}$ can be expressed as $\bm{w}_{0}^{\rm A}=\sum_{i=1}^{\operatorname{rank}(\bm{G})}\beta_{i}\bm{v}_{i}+\bar{\bm{v}}_{{\rm t},0}$, with $\bar{\bm{v}}_{{\rm t},0}$ belonging to the null space of $\bm{G}$. As the number of transmission rounds $\ell$ increase, the vectors $\bm{y}_{\ell}^{\rm A}$ and $\bm{y}_{\ell}^{\rm B}$ in (12) would be dominated by the top singular vectors $\bm{v}_{1}$ and $\bm{u}_{1}$, respectively, with a linear convergence rate [7]. Furthermore, to meet the transmit power constraint, the unit 2-norm normalization is performed in each iteration, i.e., $\bm{w}_{\ell}^{\rm B}={\bm{y}_{\ell}^{\rm B}}/{\|\bm{y}_{\ell}^{\rm B}\|_{2}}$ and $\bm{w}_{\ell+1}^{\rm A}={\bm{y}_{\ell}^{\rm A}}/{\|\bm{y}_{\ell}^{\rm A}\|_{2}}$, which does not alter the directions of $\bm{y}_{\ell}^{\rm A}$ and $\bm{y}_{\ell}^{\rm B}$. The rapid convergence rate of the power iteration scheme implies a significant pilot overhead reduction as compared to estimating the entire channel matrix.
@@ III-B Proposed Active Sensing Approach
Similarly, agent A uses its hidden state vector $\bm{h}_{L,i}^{\rm A}$ and received pilots $\bm{Y}_{L-1}^{\rm A}$ to design its precoding matrix $\bm{W}_{\rm t}$ following the same procedure as in (20), but replaces the $\operatorname{DNN}^{\rm r}$ and $\bm{Y}_{L-1}^{\rm B}$ with $\operatorname{DNN}^{\rm t}$ and $\bm{Y}_{L-1}^{\rm A}$, respectively.
The proposed active sensing framework is trained by sequentially concatenating $L$ active sensing units to learn the global policy for the entire sensing trajectory during the training stage. The loss function is set as $-\mathbb{E}\left[\log|\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})|^{2}\right]$. Furthermore, to train a neural network to be generalizable to different number of ping-pong rounds $L$, we can set the loss function as $-\mathbb{E}\left[\sum_{\ell=0}^{L-1}\log|\det(\bm{W}_{{\rm r},\ell}^{\sf H}\bm{G}\bm{W}_{{\rm t},\ell})|^{2}\right]$, where $\bm{W}_{{\rm r},\ell}$ and $\bm{W}_{{\rm t},\ell}$ are the data transmission precoding and combining matrices generated by the neural network in round $\ell$.
@@ IV Performance Evaluation for Fully Digital MIMO Systems
In this section, we evaluate the performance of the proposed active sensing framework for fully digital MIMO systems. The dimensions of the hidden states of the GRUs are set to be $N_{\rm h}^{\rm A}=N_{\rm h}^{\rm B}=512$. All the DNNs are two-layer fully connected neural networks with the dimension $[512,1024,2M_{\rm t}]$ and $\operatorname{relu}(\cdot)$ as the activation function in the hidden layers. The entire neural network involving $L$ concatenated active sensing units (as shown in Fig. 2) is implemented on PyTorch [30], which can automatically compute the gradient of the loss function with respect to the trainable parameters. We train the neural network offline using the Adam optimizer [31], with a default configuration in PyTorch and with an initial learning rate progressively decreasing from $10^{-3}$ to $10^{-5}$. In each training iteration, the batch size of training data is $1024$. We generate as much training data as needed and stop training if the performance on the validation dataset does not improve after several epochs. The testing dataset consists of $1000$ randomly generated channel matrices. In the training stage, we use the loss function $-\mathbb{E}\left[\sum_{\ell=0}^{L-1}\log|\det(\bm{W}_{{\rm r},\ell}^{\sf H}\bm{G}\bm{W}_{{\rm t},\ell})|^{2}\right]$ to train a single model that is generalizable to different number of ping-pong rounds.
@@
Figure 3: Performance comparison under the Rayleigh fading channel model for a fully digital MIMO system with $M_{\rm t}=M_{\rm r}=64$.
@@
Figure 5: Performance comparison under a ray-tracing channel model for a fully digital MIMO system with $M_{\rm t}=64,M_{\rm r}=16$ and $N_{\rm s}=2$.
@@ IV-A Rayleigh Fading Channel
In Fig. 3, we compare the performances of different methods in terms of the average objective value for varying SNRs and pilot lengths. The results of the scenario with $N_{\rm s}=2$ and $N_{\rm s}=4$ data streams are shown in Fig. 3a3d and Fig. 3e3h, respectively. In both scenarios, the neural network is trained for the case where the number of ping-pong rounds is $L=8$ (corresponding to pilot length $322LN_{\rm s}=$ 32 for $N_{\rm s}=2$ and $642LN_{\rm s}=$ 64 for $N_{\rm s}=4$).
@@
In Fig. 3e3h, we further compare the performances of different methods for the scenario $N_{\rm s}=4$. It is observed that the proposed active sensing method still achieves the best performance. However, by comparing Fig. 3c and Fig. 3g, we can see that the conventional power iteration method and summed power method is not as competitive as the channel estimation based approach when the number of data streams increases. We should also note that the conventional channel estimation based approach requires an additional error-free feedback channel to transmit the designed precoding matrix from agent B to A. The feedback channel may not be easy to realize when the SNR is low. The other approaches do not need the feedback channel since the two agents can implicitly coordinate at the ping-pong pilot transmission stage.
In Fig. 4, we further present the theoretical achievable rate (without employing water-filling) with $N_{\rm s}=2$, $M_{\rm r}=64$ and $M_{\rm t}=64$, where the data transmission SNR is set at $10$dB. From Fig. 4, it is evident that the proposed active sensing method outperforms other approaches across various pilot SNRs. This figure shows that optimizing the proposed objective function $\log|\det(\bm{W}_{\rm r}^{\sf H}\bm{G}\bm{W}_{\rm t})|^{2}$ also approximately maximizes the achievable rate.
@@ V-A System Model
Unlike the fully digital scenario, the problem of finding the optimal hybrid beamforming matrices does not have an analytical solution and is more challenging even when the channel matrix $\bm{G}$ is perfectly known. For the fully digital MIMO system, the optimal overall precoding matrix $\bm{W}_{\rm t}^{\ast}$ and overall combining matrix $\bm{W}_{\rm r}^{\ast}$ can be found as in (5), but it is not clear whether a further matrix decomposition step on $\bm{W}_{\rm t}^{\ast}$ and $\bm{W}_{\rm r}^{\ast}$ can be performed to obtain the optimal analog and digital beamformers. One approach for finding the hybrid precoding beamformers is to formulate the following matrix decomposition problem [33]: